# Charged-Current Elastic Scattering at the Electron-Ion Collider

### v1

@@ Abstract
We discuss the measurement of the charged-current elastic scattering process $e^{-}p\rightarrow\nu_{e}n$ at the Electron-Ion Collider (EIC). This process provides sensitivity to the poorly constrained axial form factor of the nucleon, which encodes the spatial distribution of weak charge. Collisions of electrons with polarized protons enable measuring the axial form factor via the $e^{-\!}\,\vec{p}\to\nu_{e}\,n$ target-spin asymmetry for the first time. We conclude that a measurement of charged-current elastic scattering at the EIC will prove very challenging.
@@ I Introduction
Understanding the sizes of nucleons and nuclei has been a primary goal of nuclear physics for over a century. The momentum-transfer dependence of elastic scattering processes provides a straightforward path to define the spatial extent of the target, and different elastic processes can provide complementary information about the distributions of quantum numbers, such as electric charge, mass, and even spin Hofstadter (1956); Polyakov and Schweitzer (2018); Miller (2019); Chen et al. (2024); Kaiser and Weise (2024); Goharipour et al. (2025).
The elastic process we will focus on in this paper is the charged-current elastic (CCE) scattering process, $e^{-}p\rightarrow\nu_{e}n$ (Fig. 1), whereby the structure of the proton is observed through the lens of a $W^{-}$. The fact that in this process the $W^{-}$ nominally couples only to left-handed up-type quarks in the proton means that it serves as a unique filter for studying flavor- and spin-dependent contributions to the nucleon structure. The form factor associated with this process is the axial form factor [^1] Also called the nucleon axial-vector form factor; some literature also uses “nucleon transition axial form factor”, since technically it describes a transition from $n\rightarrow p$ or vice versa., $F_{A}(t=-Q^{2})$ [^2] In elastic scattering, the momentum transfer to the proton and the lepton are equivalent. Therefore, $|t|$ and $Q^{2}$ can be used interchangeably, and we use only $|t|$ for convenience in the following text. Llewellyn Smith (1972). This form factor parametrizes the spatial distribution of axial currents in the nucleon. The so-called nucleon “axial radius”, $r_{A}$, is defined as
@@
where $F_{A}^{\prime}(0)$ and $F_{A}(0)$ refer to the slope and the value of the form factor at zero momentum transfer, respectively Chen et al. (2024). The authors of Ref. Chen et al. (2024) highlight $r_{A}$ as the primary component of a proton “spin radius”, defined as $\langle r_{\mathrm{spin}}^{2}\rangle=r_{A}^{2}+\frac{1}{4M^{2}}(1+\frac{2F_{p}(0)}{F_{A}(0)})$, where $F_{p}(0)$ is the induced pseudoscalar form factor at zero momentum transfer. $r_{A}$ provides novel information about how spin is arranged in the nucleon at a global scale, and measurement of $r_{A}$ is an attractive goal for future experiments.
The value of $F_{A}(0)$ is known to high precision from $\beta$-decay processes. The measured value is $F_{A}(0)=-1.2723\pm 0.0023$, which sets the normalization of the form factor. The axial form factor is also a common target of lattice calculations Alexandrou et al. (2017, 2021); Capitani et al. (2019); Gupta et al. (2017); Bali et al. (2020); Tsuji et al. (2025). For the purpose of extracting the axial radius of the nucleon, the necessary quantity to measure is the slope of the form factor as $t\rightarrow 0$, denoted here as $F^{\prime}_{A}(t\rightarrow 0)$.
@@
The landscape of existing experimental measurements of $F_{A}$ includes quasi-elastic neutrino scattering on nuclei, pion electroproduction on protons, and a measurement of $\bar{\nu_{\mu}}p\rightarrow\mu^{+}n$ in the organic scintillator of the MINERvA experiment Choi et al. (1993); Liesenfeld et al. (1999); Abe et al. (2020); Aguilar-Arevalo et al. (2010); Cai et al. (2023). Both the neutrino data with nuclear targets and the pion electroproduction data suffer from large epistemic uncertainties and model dependence Alvarez-Ruso et al. (2018); Haberzettl (2000). The MINERvA data therefore provide the cleanest existing measurement of the axial form factor, and the value of $r_{A}$ extracted from that data is $0.73\pm 0.17$ fm, which is consistent with the charge radius of the proton within the uncertainties.
The goal of this manuscript is to investigate the feasibility of measuring the CCE process at the EIC. The highly polarized proton beam at the EIC opens the door for a totally novel measurement of $F_{A}$ via the target-spin asymmetry of the CCE process. A polarized proton target is infeasible in neutrino scattering experiments due to the large target mass required to acquire adequate statistics for a measurement. Ref. Tomalak (2021) calculates the target-spin asymmetries for the crossed process $\nu n\rightarrow l^{-}p$ for both longitudinal and transverse target polarizations. These target-spin asymmetries are calculated to be very large, nearly 100% at low $Q^{2}$, and also highly sensitive to the axial form factor. A combined measurement of the target-spin asymmetry and the $|t|$-dependence of the cross section of the CCE process in $ep$ scattering therefore provides a unique way to access $F_{A}$ that is complementary to neutrino scattering.
@@ II Measurement of Charged-Current Elastic Scattering at the EIC
Compared to existing and proposed measurements of the axial form factors via (anti)neutrino scattering, performing the measurement with an electron beam provides some obvious benefits. First, the beam energy is precisely known. Electron beams can also generally attain much greater intensities than neutrino beams. Another useful feature is that electron beams often have longitudinal polarization. Since the CCE process is 100% parity-violating, only the electron polarization state where the electron spin is anti-aligned with its momentum can participate in the signal reaction. Selecting the aligned electron polarization state can therefore provide a data-driven estimate of the background. These aspects, among others, have also been discussed in a recent proposal to measure the axial form factor in electron scattering at Jefferson Lab Averett et al. (2025).
@@
Figure 2: Compilation of data for the reaction $\nu_{\mu}n\rightarrow\mu^{-}p$ reproduced from Ref. Formaggio and Zeller (2012). By crossing symmetry, this cross section should be identical to $e^{-}p\rightarrow\nu_{e}n$, neglecting lepton-mass effects which vanish at higher energies.
The EIC, at its design luminosity, will deliver around 100 fb-1 per year in the 10 GeV electron on 275 GeV proton (henceforth denoted 10x275) configuration Abdul Khalek et al. (2022). The corresponding energy of the electron in the frame where the proton is at rest for the 10x275 GeV configuration is 5.9 TeV. To estimate the value of the CCE cross section at the EIC we perform a linear fit of the world data on $\nu_{\mu}n\rightarrow\mu^{-}p$ shown in Fig. 2 for $E_{\nu}>3$ GeV. Utilizing crossing symmetry and neglecting differences between muons and electrons, we arrive at an expected cross section for CCE at $E_{\mathrm{lab}}=5.9$ TeV of roughly $7\times 10^{-39}~\mathrm{cm}^{2}$, i.e. 7 femtobarns. Since this cross section corresponds only to left-handed leptons and the EIC beam will be an equal mix of left- and right-handed electrons, the expected event rate for the EIC at maximum luminosity is around 350 events per year. CCE is therefore a prime example of a process that would have been impossible to measure at HERA but is potentially within reach thanks to the high luminosity of the EIC.
Currently, the EIC project includes one detector, the Electron-Proton/Ion Collider (ePIC) experiment, with the possibility of a second detector. The benefit of searching for the CCE process at the EIC is that the produced neutrons will be at very high energies, making them easier to detect and distinguish from non-physics backgrounds. The detector subsystem of ePIC that will measure the elastically scattered neutron is the Zero-Degree Calorimeter (ZDC), which employs an SiPM-on-tile design Blazey et al. (2009); Simon and Soldner (2010); Liu et al. (2015); Magnan (2017); Sefkow and Simon (2019). This design is expected to achieve excellent resolutions on both neutron energies and neutron positions Milton et al. (2025); Zhang et al. (2025). The anticipated energy resolution using a graph neural network reconstruction technique is $\frac{\sigma_{E}}{E}=29\%/\sqrt{E}\oplus 2\%$, and the position resolution for neutrons near the beam energy is on the order of 0.04 mrad. As can be seen from Fig. 3, for $|t|<2\mathrm{~GeV^{2}}$ the final-state neutron has a momentum indistinguishable from the beam momentum. Using an angular resolution of 0.04 mrad, the resolution on $|t|$ for the CCE process is around 5%. However, the angular divergence of the proton beam in the highest luminosity configuration actually dominates the resolution, providing a contribution to the angular resolution of 0.119 mrad Electron-Ion Collider Project (2023).
@@ III Background
Since the CCE process is reconstructed from a single neutron, processes which produce high-energy neutrons serve as a background. To understand the background from inclusive leading neutron photoproduction, $\gamma+p\rightarrow n+X$, the ZEUS data at $\sqrt{s}=300$ and $\sqrt{s}=318$ GeV provide a useful starting point Chekanov et al. (2002, 2007). ZEUS provides photoproduction cross sections as a function of $x_{L}$, defined as
@@
The result is that the signal-to-background for CCE events at the EIC is expected to be around $10^{-7}$ to $10^{-8}$. Unfortunately, the shape of the $|t|$-distribution of the background will likely be similar to that of the signal, although it can be measured in several ways. The shape of the background can be measured using the right-handed electron beam bunches as a control sample or by using events where a scattered electron is detected in the Low-$Q^{2}$ tagger. To have a measurement of $5\sigma$ statistical significance for the expected 350 events in one year of running, only around 5000 background events can survive the vetoing process, assuming a perfect knowledge of the background distribution. This means that the required background rejection factor necessary for 100 fb-1 is on the order of $10^{7}$, decreasing to $10^{6}$ if 1 ab-1 is collected. Rejection of background at these levels is common in the realm of rare decay experiments but is certainly a unique challenge in a collider environment. At colliders, efficiency for reconstruction of particles is typically valued over efficiency for vetoing particles, and the latter is in general poorly studied for the planned EIC detector subsystems. For most exclusive channels, which would benefit the most from high veto efficiency, the kinematic constraints imposed by exclusivity are enough to reduce backgrounds to the required levels. Since the effectiveness of the ePIC subdetectors for vetoing has not been studied in detail (with the exception of Ref. Aschenauer et al. (2025)), a quantitative estimate of how much the background can be reduced is at present beyond the scope of this work. The remainder of this section will be dedicated to some strategies that could be used to reduce the background for the CCE process.
@@
While the angular resolution from the beam divergence dominates the measurement of $|t|$, the energy resolution of the ZDC still plays a vital role in the measurement, namely for vetoing background from lower $x_{L}$ leading neutron events. No inelastic process can produce neutrons as close to the beam energy as the CCE process. Therefore the ability to reject neutrons with incident energies below the endpoint will drastically reduce the background. At the endpoint energy, the expected energy resolution of the ZDC is 2.6%, i.e. 7.2 GeV. Crucially, this energy resolution does not depend on $\theta_{n}$ for the range of scattering angles corresponding to $|t|$ ¡ 2 GeV2. It can be seen from Fig. 3 that all CCE data reside at $x_{L}=1$, to a good approximation. A cut on neutron energies measured more than $2\sigma_{E}$ away from the endpoint $x_{L}=1$ would remove $x_{L}<0.948$. In Ref. Chekanov et al. (2007), ZEUS measures that in a bin of $0.95<x_{L}<1.00$, the leading neutron cross section is around 7 times smaller than the cross section in the bin from $0.90<x_{L}<0.95$. Since we are concerned only with the kinematic endpoint of $x_{L}\approx 1$, we can gain around a factor of 10 suppression compared to the $x_{L}>0.88$ cross section by selecting neutrons within $2\sigma_{E}$ of the beam energy. A better energy resolution on the ZDC, albeit hard to imagine since the assumed resolution is already better than any previous hadronic calorimeter at a collider, would help by reducing the region of neutron energies that need to be accepted.
@@
The most straightforward route to veto non-CCE events is to veto any event containing a scattered electron, including those associated with photoproduction events. Electroproduction of leading neutrons for $Q^{2}>1$ GeV2 should be able to be efficiently vetoed by measuring the scattered electron in the ePIC central detector. Based on their anticipated efficiencies, the lead tungstate backward calorimeter Houžvička et al. (2022); Philip et al. (2024); Mkrtchyan et al. (2025), proximity-focusing RICH detector Bhattacharya et al. (2023); Chatterjee (2025), and backward tracker Sidoretti et al. (2025a, b); Gonella (2024); Li (2024) should combine to a background rejection factor similar to the one required for $Q^{2}>1$ GeV2. Some of the $Q^{2}<1$ GeV2 phase space will be covered by the far-backward detectors Gardner et al. (2023). In principle, the ePIC low-$Q^{2}$ tagger has acceptance for $10^{-7}<Q^{2}<0.01$ GeV2, although this detector cannot be used as a reliable veto since it has a relatively low acceptance and a high occupancy per bunch crossing from Bethe-Heitler events Gardner et al. (2023).
@@
It is important to note that all background reactions must have a charged hadron in the final state while the signal reaction does not. Vetoing these charged hadrons provides essentially the only hope to reduce the background to a manageable level. Charged hadrons produced near the proton beam ($\eta>4.5$) will be swept out of the beampipe at the B0 dipole and at that point can be vetoed to reduce the background. To get an idea of the distributions of hadrons in leading-neutron events, we use Pythia6 Sjostrand et al. (2006). The events were generated with no cuts on kinematic variables, i.e. including photoproduction, and the distribution of charged hadron momenta and pseudorapidity are shown in Fig. 5. The band at large $\eta$ and momentum corresponds to $\Delta^{+}$ baryon production and decay to $\pi^{+}n$. It can be seen that vetoing charged hadrons in the range of $\eta>3.5$ is crucial to veto events with high-energy neutrons produced from low-mass baryon resonances.
Figure 5: Momentum vs. pseudorapidity for the highest-momentum charged particle in an event with an $x_{L}>0.95$ leading neutron as simulated by Pythia6 Sjostrand et al. (2006).

### v2

@@ Abstract
We discuss the measurement of the charged-current elastic scattering process $e^{-}p\rightarrow\nu_{e}n$ at the Electron-Ion Collider (EIC). This process provides sensitivity to the poorly constrained axial form factor of the nucleon, which encodes the spatial distribution of weak charge. Collisions of electrons with polarized protons enable measuring the axial form factor via the $e^{-\!}\,\vec{p}\to\nu_{e}\,n$ target-spin asymmetry for the first time. We conclude that a measurement of charged-current elastic scattering at the EIC will, perhaps unsurprisingly, prove very challenging. However, with dedicated instrumentation at a second EIC detector, the measurement may be possible.
@@ I Introduction
Understanding the sizes of nucleons and nuclei has been a primary goal of nuclear physics for over a century. The momentum-transfer dependence of elastic scattering processes provides a straightforward path to define the spatial extent of the target, and different elastic processes can provide complementary information about the distributions of quantum numbers, such as electric charge, mass, and even spin Hofstadter1956; PolyakovSchweitzer2018; Miller:2019DefiningRp; Chen:2024oxx; KaiserWeise:2024Sizes; Goharipour:2025yxm.
The elastic process we will focus on in this paper is the charged-current elastic (CCE) scattering process, $e^{-}p\rightarrow\nu_{e}n$ (Fig. 1), whereby the structure of the proton is observed through the lens of a $W^{-}$. The fact that in this process the $W^{-}$ nominally couples only to left-handed up-type quarks in the proton means that it serves as a unique filter for studying flavor- and spin-dependent contributions to the nucleon structure. The form factor associated with this process is the axial form factor [^1] Also called the nucleon axial-vector form factor; some literature also uses “nucleon transition axial form factor”, since technically it describes a transition from $n\rightarrow p$ or vice versa., $F_{A}(t=-Q^{2})$ [^2] In elastic scattering, the momentum transfer to the proton and the lepton are equivalent. Therefore, $|t|$ and $Q^{2}$ can be used interchangeably, and we use only $|t|$ for convenience in the following text. LlewellynSmith1972. This form factor parametrizes the spatial distribution of axial currents in the nucleon. The so-called nucleon “axial radius”, $r_{A}$, is defined as
@@
where $F_{A}^{\prime}(0)$ and $F_{A}(0)$ refer to the slope and the value of the form factor at zero momentum transfer, respectively Chen:2024oxx. The authors of Refs. Chen:2024oxx; Chen:2024ksq highlight $r_{A}$ as the primary component of a proton “spin radius”, defined as $\langle r_{\mathrm{spin}}^{2}\rangle=r_{A}^{2}+\frac{1}{4M^{2}}(1+\frac{2F_{p}(0)}{F_{A}(0)})$, where $F_{p}(0)$ is the induced pseudoscalar form factor at zero momentum transfer. $r_{A}$ provides novel information about how spin is arranged in the nucleon at a global scale, and measurement of $r_{A}$ is an attractive goal for future experiments.
The value of $F_{A}(0)$ is known to high precision from $\beta$-decay processes. The measured value is $F_{A}(0)=-1.2723\pm 0.0023$, which sets the normalization of the form factor. The axial form factor is also a common target of lattice calculations Alexandrou:2017hac; Alexandrou:2020okk; Capitani:2017qpc; Gupta:2017dwj; RQCD:2019jai; Tsuji:2025quu. For the purpose of extracting the axial radius of the nucleon, the necessary quantity to measure is the slope of the form factor as $t\rightarrow 0$, denoted here as $F^{\prime}_{A}(t\rightarrow 0)$.
@@
The landscape of existing experimental measurements of $F_{A}$ includes quasi-elastic neutrino scattering on nuclei, pion electroproduction on protons, and a measurement of $\bar{\nu_{\mu}}p\rightarrow\mu^{+}n$ in the organic scintillator of the MINERvA experiment Choi:1993PRL_AxialPionElectroprod; Liesenfeld:1999PLB_AxialFromPeepiN; T2K:2020CC0piOC; AguilarArevalo:2010MiniBooNECCQE; MINERvA:2023avz. Both the neutrino data with nuclear targets and the pion electroproduction data suffer from large epistemic uncertainties and model dependence NuSTEC:2017hzk; Haberzettl:2000sm. The MINERvA data therefore provide the cleanest existing measurement of the axial form factor, and the value of $r_{A}$ extracted from that data is $0.73\pm 0.17$ fm, which is consistent with the charge radius of the proton within the uncertainties.
The goal of this manuscript is to investigate the feasibility of measuring the CCE process at the EIC. The highly polarized proton beam at the EIC opens the door for a totally novel measurement of $F_{A}$ via the target-spin asymmetries of the CCE process. A polarized proton target is infeasible in neutrino scattering experiments due to the large target mass required to acquire adequate statistics for a measurement. Refs. Tomalak:2020zlv; Graczyk:2019xwg calculate the target-spin asymmetries for the crossed process $\nu n\rightarrow l^{-}p$ for both longitudinal and transverse target polarizations. These target-spin asymmetries are calculated to be very large, the longitudinal asymmetry $A_{UL}$ reaches nearly 100% at low $Q^{2}$ and the transverse asymmetry $A_{UT}$ is greater than 50% for $Q^{2}>0.5\mathrm{~GeV}^{2}$. Ref. Tomalak:2020zlv further demonstrates the sensitivity of the target-spin asymmetry to the axial form factor. A combined measurement of the target-spin asymmetry and the $|t|$-dependence of the cross section of the CCE process in $ep$ scattering therefore provides a unique way to access $F_{A}$ that is complementary to neutrino scattering.
@@ II Measurement of Charged-Current Elastic Scattering at the EIC
Compared to existing and proposed measurements of the axial form factors via (anti)neutrino scattering, performing the measurement with an electron beam provides some obvious benefits. First, the beam energy is precisely known. Electron beams can also generally attain much greater intensities than neutrino beams. Another useful feature is that electron beams often have longitudinal polarization. Since the CCE process is 100% parity-violating, only the electron polarization state where the electron spin is anti-aligned with its momentum can participate in the signal reaction. Selecting the aligned electron polarization state can therefore provide a data-driven estimate of the background. These aspects, among others, have also been discussed in a recent proposal to measure the axial form factor in electron scattering at Jefferson Lab JLab-PR12-25-009.
@@
Figure 2: Compilation of data for the reaction $\nu_{\mu}n\rightarrow\mu^{-}p$ reproduced from Ref. Formaggio:2012cpf. By crossing symmetry, this cross section should be identical to $e^{-}p\rightarrow\nu_{e}n$, neglecting lepton-mass effects which vanish at higher energies.
The EIC, at its design luminosity, will deliver around 100 fb-1 per year in the 10 GeV electron on 275 GeV proton (henceforth denoted 10x275) configuration AbdulKhalek:2021gbh. The corresponding energy of the electron in the frame where the proton is at rest for the 10x275 GeV configuration is 5.9 TeV. To estimate the value of the CCE cross section at the EIC we perform a linear fit of the world data on $\nu_{\mu}n\rightarrow\mu^{-}p$ shown in Fig. 2 for $E_{\nu}>3$ GeV. Utilizing crossing symmetry and neglecting differences between muons and electrons, we arrive at an expected cross section for CCE at $E_{\mathrm{lab}}=5.9$ TeV of roughly $7\times 10^{-39}~\mathrm{cm}^{2}$, i.e. 7 femtobarns. Since this cross section corresponds only to left-handed leptons and the EIC beam will be an equal mix of left- and right-handed electrons, the expected event rate for the EIC at maximum luminosity is around 350 events per year. CCE is therefore a prime example of a process that would have been impossible to measure at HERA but is potentially within reach thanks to the high luminosity of the EIC.
Currently, the EIC project includes one detector, the Electron-Proton/Ion Collider (ePIC) experiment, with the possibility of a second detector. The benefit of searching for the CCE process at the EIC is that the produced neutrons will be at very high energies, making them easier to detect and distinguish from non-physics backgrounds. The detector subsystem of ePIC that will measure the elastically scattered neutron is the Zero-Degree Calorimeter (ZDC), which employs an SiPM-on-tile design Blazey:2009zz; Simon:2010hf; Liu:2015cpe; Magnan:2017exp; Sefkow:2018rhp. This design is expected to achieve excellent resolutions on both neutron energies and neutron positions Milton:2024bqv; Zhang:2025ijg. The anticipated energy resolution using a graph neural network reconstruction technique is $\frac{\sigma_{E}}{E}=29\%/\sqrt{E}\oplus 2\%$, and the position resolution for neutrons near the beam energy is on the order of 0.04 mrad. As can be seen from Fig. 3, for $|t|<2\mathrm{~GeV^{2}}$ the final-state neutron has a momentum indistinguishable from the beam momentum. Using an angular resolution of 0.04 mrad, the resolution on $|t|$ for the CCE process is around 5%. However, the angular divergence of the proton beam in the highest luminosity configuration actually dominates the resolution, providing a contribution to the angular resolution of 0.119 mrad EIC_MasterParameter_Table_2023.
@@ III Background
Since the CCE process is reconstructed from a single neutron, processes which produce high-energy neutrons serve as a background. To understand the background from inclusive leading neutron photoproduction, $\gamma+p\rightarrow n+X$, the ZEUS data at $\sqrt{s}=300$ and $\sqrt{s}=318$ GeV provide a useful starting point ZEUS:2002gig; ZEUS:2007knd. ZEUS provides photoproduction cross sections as a function of $x_{L}$, defined as
@@
The result is that the signal-to-background for CCE events at the EIC is expected to be around $10^{-7}$ to $10^{-8}$. Unfortunately, the shape of the $|t|$-distribution of the background will likely be similar to that of the signal, although it can be measured in several ways. The shape of the background can be measured using the right-handed electron beam bunches as a control sample or by using events where a scattered electron is detected in the Low-$Q^{2}$ tagger. To have a measurement of $5\sigma$ statistical significance for the expected 350 events in one year of running, only around 5000 background events can survive the vetoing process, assuming a perfect knowledge of the background distribution. This means that the required background rejection factor necessary for 100 fb-1 is on the order of $10^{7}$, decreasing to $10^{6}$ if 1 ab-1 is collected. Rejection of background at these levels is common in the realm of rare decay experiments but is certainly a unique challenge in a collider environment. At colliders, efficiency for reconstruction of particles is typically valued over efficiency for vetoing particles, and the latter is in general poorly studied for the planned EIC detector subsystems. For most exclusive channels, which would benefit the most from high veto efficiency, the kinematic constraints imposed by exclusivity are enough to reduce backgrounds to the required levels. Since the effectiveness of the ePIC subdetectors for vetoing has not been studied in detail (with the exception of Ref. Aschenauer:2025mku), a quantitative estimate of how much the background can be reduced is at present beyond the scope of this work. The remainder of this section will be dedicated to some strategies that could be used to reduce the background for the CCE process.
@@
While the angular resolution from the beam divergence dominates the measurement of $|t|$, the energy resolution of the ZDC still plays a vital role in the measurement, namely for vetoing background from lower $x_{L}$ leading neutron events. No inelastic process can produce neutrons as close to the beam energy as the CCE process. Therefore the ability to reject neutrons with incident energies below the endpoint will drastically reduce the background. At the endpoint energy, the expected energy resolution of the ZDC is 2.6%, i.e. 7.2 GeV. Crucially, this energy resolution does not depend on $\theta_{n}$ for the range of scattering angles corresponding to $|t|<2~\mathrm{GeV}^{2}$. It can be seen from Fig. 3 that all CCE data reside at $x_{L}=1$, to a good approximation. A cut on neutron energies measured more than $2\sigma_{E}$ away from the endpoint $x_{L}=1$ would remove $x_{L}<0.948$. In Ref. ZEUS:2007knd, ZEUS measures that in a bin of $0.95<x_{L}<1.00$, the leading neutron cross section is around 7 times smaller than the cross section in the bin from $0.90<x_{L}<0.95$. Since we are concerned only with the kinematic endpoint of $x_{L}\approx 1$, we can gain around a factor of 10 suppression compared to the $x_{L}>0.88$ cross section by selecting neutrons within $2\sigma_{E}$ of the beam energy. A better energy resolution on the ZDC, albeit hard to imagine since the assumed resolution is already better than any previous hadronic calorimeter at a collider, would help by reducing the region of neutron energies that need to be accepted.
@@
The most straightforward route to veto non-CCE events is to veto any event containing a scattered electron, including those associated with photoproduction events. Electroproduction of leading neutrons for $Q^{2}>1$ GeV2 should be able to be efficiently vetoed by measuring the scattered electron in the ePIC central detector. Based on their anticipated efficiencies, the lead tungstate backward calorimeter Houzvicka:2022pny; Philip:2024opz; Mkrtchyan:2025euu, proximity-focusing RICH detector Bhattacharya:2023ddr; Chatterjee:2024zrn, and backward tracker Sidoretti:2025tja; Sidoretti:2025xeh; Gonella:2024cxk; Li:2023hjr should combine to a background rejection factor similar to the one required for $Q^{2}>1$ GeV2. Some of the $Q^{2}<1$ GeV2 phase space will be covered by the far-backward detectors Gardner:2023lly. In principle, the ePIC low-$Q^{2}$ tagger has acceptance for $10^{-7}<Q^{2}<0.01$ GeV2, although this detector cannot be used as a reliable veto since it has a relatively low acceptance and a high occupancy per bunch crossing from Bethe-Heitler events Gardner:2023lly.
@@
It is important to note that all background reactions must have a charged hadron in the final state while the signal reaction does not. Vetoing these charged hadrons provides essentially the only hope to reduce the background to a manageable level. Charged hadrons produced near the proton beam ($\eta>4.5$) will be swept out of the beampipe at the B0 dipole and at that point can be vetoed to reduce the background. To get an idea of the distributions of hadrons in leading-neutron events, we use Pythia6 Sjostrand:2006za. The events were generated with no cuts on kinematic variables, i.e. including photoproduction, and the distribution of charged hadron momenta and pseudorapidity are shown in Fig. 5. The band at large $\eta$ and momentum corresponds to $\Delta^{+}$ baryon production and decay to $\pi^{+}n$. It can be seen that vetoing charged hadrons in the range of $\eta>3.5$ is crucial to veto events with high-energy neutrons produced from low-mass baryon resonances.
Figure 5: Momentum vs. pseudorapidity for the highest-momentum charged particle in an event with an $x_{L}>0.95$ leading neutron as simulated by Pythia6 Sjostrand:2006za.